# Theoretical Limits of Protocols for Distinguishing Different Unravelings

### v1

@@ I Introduction
Stochastic Schrödinger equations (SSEs) appear in a variety of contexts Wiseman and Milburn (2009, 1993); Carmichael (2009); Dalibard et al. (1992); Nielsen and Chuang (2010); Preskill (2018); Breuer and Petruccione (2002); Gisin and Percival (1992); Mølmer et al. (1993); Daley (2014); Diósi (1998); Ghirardi et al. (1990); Kafri et al. (2014); Tilloy and Diósi (2016, 2017); Weiss (2012); Hänggi et al. (1990); Wiseman and Diósi (2001); Caiaffa et al. (2017); Jacobs (2014); Wiseman (1996); Carmichael (2007); Piccitto et al. (2022); Cao et al. (2019); Vovk and Pichler (2024); Piccitto et al. (2024); Bassi and Ippoliti (2006); Genoni et al. (2016); Wu et al. (2024); Piñol et al. (2024); Albarelli et al. (2017); Magrini et al. (2021); Zhang and Mølmer (2017); Belenchia et al. (2020); Zhang et al. (2017). In quantum optics, they are used to describe the dynamics of systems under continuous measurement Wiseman and Milburn (2009, 1993), for different photon detection schemes Carmichael (2009); Dalibard et al. (1992). In quantum computation, they are employed to analyze the impact of noise and errors on qubits, in light of fault-tolerant quantum computing Nielsen and Chuang (2010); Preskill (2018). Within the theory of open quantum systems, they allow to describe the reduced system dynamics at the wave function level Breuer and Petruccione (2002); Gisin and Percival (1992). Numerical simulations based on SSEs, including quantum Monte Carlo methods, enable the simulation of larger quantum systems, making SSEs a common tool for tackling computationally intensive problems Mølmer et al. (1993); Daley (2014). Stochastic Schrödinger equations are also used in the foundations of quantum mechanics to model the spontaneous collapse of the wave function Diósi (1998); Ghirardi et al. (1990); more recently, they have been use to model the coupling between classical gravity and quantum matter Kafri et al. (2014); Tilloy and Diósi (2016, 2017) offering an alternative to quantum gravity.
@@
The association between master equations and stochastic unravelings is not unique: in general, there are infinite ways to unravel a master equation via SSEs Wiseman and Diósi (2001); Caiaffa et al. (2017), as also shown later on. For example, in quantum optics, the detection of photons at random times leaking out of an optical cavity is described by a stochastic equation for the state of the system, with the randomness encoded in a Poisson increment Jacobs (2010, 2014). In homodyne detection, a photodetector is used to perform a Gaussian continuous measurement of an optical mode, where instead of a direct detection, light is first interfered (with light of the same frequency) in a beam splitter. The resulting homodyne stochastic equation is defined in terms of a Wiener process Wiseman (1996); Carmichael (2007); Jacobs (2014). These are two different measurement processes; yet at the statistical level, when averaged over the noise, they are indistinguishable from each other Jacobs (2014).
@@
Recently, it has been suggested Piñol et al. (2024) that these nonlinear quantities can be used to operationally distinguish different unravelings of a given master equation. This question is both interesting and relevant to the theory of open quantum systems. However, here we show that this is not the case.
The reason is that nonlinear quantities are either reconstructed from the unraveling associated with the measurement procedure, and thus quantify differences that are embodied in the choice of the unraveling—i.e., the operational distinction is already made in selecting the measurement procedure, not in computing the associated nonlinear quantities—or they are not accessible at all. The latter point is particularly crucial, as their accessibility would enable the development of a protocol for superluminal signaling, which is (supposed to be) forbidden.
@@ II Nonlinear and linear unravelings of the master equation
where $\langle\hat{L}\rangle_{t}=\bra{\psi_{t}}\hat{L}\ket{\psi_{t}}$, $W_{t}$ is a standard Wiener process and $\xi=\xi^{\text{\tiny{R}}}+i\xi^{\text{\tiny{I}}}$ is a complex parameter such that $\xi^{\text{\tiny{R}}}\geq 0$ and $|\xi|=1$. Throughout this work we will use the Itô formalism. With the exception of the limiting case $\xi^{\text{\tiny R}}=0$, one can show  [^2] See Appendix A for the details that the above family of unravelings corresponds to different types of Gaussian continuous measurement Jacobs (2014), with measurement record
@@ II.1 Free-particle
is solution for both unravelings, where $a_{t}$ and $\gamma_{t}$ are complex functions of time, whereas $\overline{x}_{t}$ and $\overline{k}_{t}$ are real. It can be shown  [^4] See Appendix B, where we present the explicit time dependence of $a_{t}$, $\gamma_{t}$, $\overline{x}_{t}$, $\overline{k}_{t}$, and we derive the free particle dynamics in the different unravelings that for the nonlinear unraveling of Eq. (4), with the choices in Eq. (8) for $\hat{H}$ and $\hat{L}$, we have:
@@
with $b,c$ and $k$ constants 555cf. Eq. (38). On the other hand, for the linear unraveling, we obtain:
@@
Both quantities differ from the variance $\text{Var}_{t}(\hat{x})=\mathbb{E}_{\omega}[\langle\hat{x}^{2}\rangle_{t}]-(\mathbb{E}_{\omega}[\langle\hat{x}\rangle])^{2}_{t}=\text{Tr}[\hat{x}^{2}\hat{\rho}_{t}]-(\text{Tr}[\hat{x}\hat{\rho}_{t}])^{2}$  [^6] In general, we have $\mathbb{E}_{\omega}[\text{Var}(\hat{O}_{i},\hat{O}_{j})]=\frac{1}{2}\mathbb{E}_{\omega}\langle\{\hat{O}_{i},\hat{O}_{j}\}\rangle_{t}-\mathbb{E}_{\omega}[\langle\hat{O}_{i}\rangle_{t}]\mathbb{E}_{\omega}[\langle\hat{O}_{j}\rangle_{t}]$, obtained from the density matrix, which by construction is unraveling-independent. Straightforward calculations show that
@@ II.2 Spin-$1/2$ particle
with $\hat{\sigma}_{z}$ the Pauli matrix and $\nu$ a coupling frequency. In this case, the solution of the stochastic equations for both the nonlinear and linear unraveling can be written explicitly. For the nonlinear unraveling given by Eq. (4), the solution is  [^7] See Appendix D, where we derive an explicit solution for the Spin-$1/2$ particle in the nonlinear unraveling:
@@
On the more formal level, following the same argument as in in Ref. Adler et al. (2001) one can show that  [^8] See Appendix  E, where we show the collapse of the wavefunction in the nonlinear unraveling:
@@
We now turn to the question of whether the nonlinear quantities introduced earlier can be operationally used to distinguish between different unravelings. From the analysis in the previous section, as well as existing literature Wiseman and Diósi (2001); Caiaffa et al. (2017), it is evident that these quantities differentiate between unravelings of the same master equation. However, this distinction has so far been purely mathematical. The key question, therefore, is whether they also provide a means to experimentally distinguish between unravelings—i.e., whether they correspond to implementable procedures beyond their mathematical definition.
@@ IV Conclusion
From a mathematical standpoint, it is well known that unravelings can be distinguished Wiseman and Diósi (2001); Caiaffa et al. (2017), not least because the very notion of different unravelings implies some form of differentiation. Thus, the real question is whether they can be distinguished operationally when they correspond to different measurement procedures.

### v2

@@ I Introduction
Stochastic Schrödinger equations (SSEs) appear in a variety of contexts Wiseman and Milburn (2009, 1993); Carmichael (2009); Dalibard et al. (1992); Nielsen and Chuang (2010); Preskill (2018); Breuer and Petruccione (2002); Gisin and Percival (1992); Mølmer et al. (1993); Daley (2014); Diósi (1998); Ghirardi et al. (1990a); Kafri et al. (2014); Tilloy and Diósi (2016, 2017); Weiss (2012); Hänggi et al. (1990); Ghirardi et al. (1990b); Pearle (1997); Wiseman and Diósi (2001); Adler and Bassi (2007); Caiaffa et al. (2017); Jacobs (2014); Wiseman (1996); Carmichael (2007); Piccitto et al. (2022); Cao et al. (2019); Vovk and Pichler (2024); Piccitto et al. (2024); Bassi and Ippoliti (2006); Genoni et al. (2016); Wu et al. (2024); Piñol et al. (2024); Albarelli et al. (2017); Magrini et al. (2021); Zhang and Mølmer (2017); Belenchia et al. (2020); Zhang et al. (2017). In quantum optics, they are used to describe the dynamics of systems under continuous measurement Wiseman and Milburn (2009, 1993), for different photon detection schemes Carmichael (2009); Dalibard et al. (1992). In quantum computation, they are employed to analyze the impact of noise and errors on qubits, in light of fault-tolerant quantum computing Nielsen and Chuang (2010); Preskill (2018). Within the theory of open quantum systems, they allow to describe the reduced system dynamics at the wave function level Breuer and Petruccione (2002); Gisin and Percival (1992). Numerical simulations based on SSEs, including quantum Monte Carlo methods, enable the simulation of larger quantum systems, making SSEs a common tool for tackling computationally intensive problems Mølmer et al. (1993); Daley (2014). Stochastic Schrödinger equations are also used in the foundations of quantum mechanics to model the spontaneous collapse of the wave function Diósi (1998); Ghirardi et al. (1990a); more recently, they have been use to model the coupling between classical gravity and quantum matter Kafri et al. (2014); Tilloy and Diósi (2016, 2017) offering an alternative to quantum gravity.
@@
The association between master equations and stochastic unravelings is not unique: in general, there are infinite ways to unravel a master equation via SSEs Ghirardi et al. (1990b); Pearle (1997); Wiseman and Diósi (2001); Adler and Bassi (2007); Caiaffa et al. (2017), as also shown later on. For example, in quantum optics, the detection of photons at random times leaking out of an optical cavity is described by a stochastic equation for the state of the system, with the randomness encoded in a Poisson increment Jacobs (2010, 2014). In homodyne detection, a photodetector is used to perform a Gaussian continuous measurement of an optical mode, where instead of a direct detection, light is first interfered (with light of the same frequency) in a beam splitter. The resulting homodyne stochastic equation is defined in terms of a Wiener process Wiseman (1996); Carmichael (2007); Jacobs (2014). These are two different measurement processes; yet at the statistical level, when averaged over the noise, they are indistinguishable from each other Jacobs (2014).
@@
Recently, it has been suggested Piñol et al. (2024) that these nonlinear quantities can be used to operationally distinguish different unravelings of a given master equation. This question is both interesting and relevant to the theory of open quantum systems.
We argue that nonlinear quantities are either reconstructed from the unraveling associated with the measurement procedure, and thus quantify differences that are embodied in the choice of the unraveling—i.e., the operational distinction is already made in selecting the measurement procedure, not in computing the associated nonlinear quantities—or they are not accessible at all. The latter point is particularly crucial, as their accessibility would enable the development of a protocol for superluminal signaling, which is (supposed to be) forbidden.
@@ II Nonlinear and linear unravelings of the master equation
where $\langle\hat{L}\rangle_{t}=\bra{\psi_{t}}\hat{L}\ket{\psi_{t}}$, $W_{t}$ is a standard Wiener process and $\xi=\xi^{\text{\tiny{R}}}+i\xi^{\text{\tiny{I}}}$ is a complex parameter such that $\xi^{\text{\tiny{R}}}\geq 0$ and $|\xi|=1$. Throughout this work we will use the Itô formalism. With the exception of the limiting case $\xi^{\text{\tiny R}}=0$, one can show  [^2] See Appendix A for the details. that the above family of unravelings corresponds to different types of Gaussian continuous measurement Jacobs (2014), with measurement record
@@ II.1 Free-particle
is solution for both unravelings, where $a_{t}$ and $\gamma_{t}$ are complex functions of time, whereas $\overline{x}_{t}$ and $\overline{k}_{t}$ are real. It can be shown  [^4] See Appendix B, where we present the explicit time dependence of $a_{t}$, $\gamma_{t}$, $\overline{x}_{t}$, $\overline{k}_{t}$, and we derive the free particle dynamics in the different unravelings. that for the nonlinear unraveling of Eq. (4), with the choices in Eq. (8) for $\hat{H}$ and $\hat{L}$, we have:
@@
with $b,c$ and $k$ constants 555cf. Eq. (38).. On the other hand, for the linear unraveling, we obtain:
@@
Both quantities differ from the variance $\text{Var}_{t}(\hat{x})=\mathbb{E}_{\omega}[\langle\hat{x}^{2}\rangle_{t}]-(\mathbb{E}_{\omega}[\langle\hat{x}\rangle])^{2}_{t}=\text{Tr}[\hat{x}^{2}\hat{\rho}_{t}]-(\text{Tr}[\hat{x}\hat{\rho}_{t}])^{2}$  [^6] In general, we have $\mathbb{E}_{\omega}[\text{Var}(\hat{O}_{i},\hat{O}_{j})]=\frac{1}{2}\mathbb{E}_{\omega}\langle\{\hat{O}_{i},\hat{O}_{j}\}\rangle_{t}-\mathbb{E}_{\omega}[\langle\hat{O}_{i}\rangle_{t}]\mathbb{E}_{\omega}[\langle\hat{O}_{j}\rangle_{t}]$., obtained from the density matrix, which by construction is unraveling-independent. Straightforward calculations show that
@@ II.2 Spin-$1/2$ particle
with $\hat{\sigma}_{z}$ the Pauli matrix and $\nu$ a coupling frequency. In this case, the solution of the stochastic equations for both the nonlinear and linear unraveling can be written explicitly. For the nonlinear unraveling given by Eq. (4), the solution is  [^7] See Appendix D, where we derive an explicit solution for the Spin-$1/2$ particle in the nonlinear unraveling.:
@@
On the more formal level, following the same argument as in in Ref. Adler et al. (2001) one can show that  [^8] See Appendix  E, where we show the collapse of the wavefunction in the nonlinear unraveling.:
@@
We now turn to the question of whether the nonlinear quantities introduced earlier can be operationally used to distinguish between different unravelings. From the analysis in the previous section, as well as existing literature Ghirardi et al. (1990b); Pearle (1997); Wiseman and Diósi (2001); Adler and Bassi (2007); Caiaffa et al. (2017), it is clear that these quantities differentiate between unravelings of the same master equation. However, this distinction has so far been purely mathematical. The key question, therefore, is whether they also provide a means to experimentally distinguish between unravelings—i.e., whether they correspond to implementable procedures beyond their mathematical definition.
@@ IV Conclusion
From a mathematical standpoint, it is well known that unravelings can be distinguished Ghirardi et al. (1990b); Pearle (1997); Wiseman and Diósi (2001); Adler and Bassi (2007); Caiaffa et al. (2017), not least because the very notion of different unravelings implies some form of differentiation. Thus, the real question is whether they can be distinguished operationally when they correspond to different measurement procedures.